# Random local access for sampling $k$-SAT solutions

### v2

@@ Appendix A Correctness of Algorithm 6
Let $\mathcal{T}$ denote the set of 2-trees of $V(G)$ of size $\eta+1$ that intersects with $W^{\prime}$. Since $|W^{\prime}|\leq d$ and $G$ has maximum degree at most $16k^{2}d-$ 1, by A.4, we have
@@
where the last step used the assumption that $\theta\geq 0.4$ and $236ed^{3}k^{4}\cdot 0.6^{\alpha k}\leq 1/$ 2. This implies that
@@
Setting $\eta=(3+c)\log n$, we get the following correctness statement on Algorithm 6.
Corollary A.11.
@@
Since we have picked $T$ in Algorithm 4 sufficiently large so that $d_{\operatorname{\mathsf{TV}}}(X_{0},\nu)\leq n^{-(2+c)}$, we get that the joint output $(\mathsf{MarginSample}(u))_{u\in\mathcal{M}}$ satisfies $d_{\operatorname{\mathsf{TV}}}((\mathsf{MarginSample}(u))_{u\in\mathcal{M}},\nu)\leq n^{-c}$, proving Proposition 4.6.
Appendix B Efficiency of Algorithm 6
We now move on to show the efficiency of $\mathsf{Glauber}(t)$ for all $t\leq 0$. Observe that for every $r\geq 48dk^{4}$, by Proposition A.5, we have
$\mathbb{P}[|R|\geq r]\leq 2^{-\frac{r}{48dk^{4}}}.$ (5)
@@ Appendix B Efficiency of Algorithm 6
We start by upper bounding the size of the final sets $V^{\prime}$ and $\Lambda$ in Algorithm 6 in terms of $|R|$.
Lemma B.1.
The $V^{\prime}$ and $\Lambda$ in Line 19 of Algorithm 6 satisfy
@@
Proof.
By Line 10 of Algorithm 6, we know that for all $u\in\Lambda$, there exists clause $C\in\mathcal{C}$ such that $u\in\operatorname{vbl}(C)$ and $\operatorname{vbl}(C)\cap V^{\prime}\neq\emptyset$. This shows that $|\Lambda|\leq kd|V^{\prime}|$. Observe that $V^{\prime}$ contains at least $|V^{\prime}|/(\Delta+1)\geq|V^{\prime}|/(2dk)$ clauses with disjoint variable sets. Since every marked variable in the clauses added to $V^{\prime}$ have gone through at least one round of LB-Sample (per Line 5 and Line 12), we get that $|R|\geq\alpha k\cdot|V^{\prime}|/(2dk)=\alpha|V^{\prime}|/(2d)$. ∎
@@
Proposition B.2 ([18, Lemma 3.10 and Appendix A]).
Let $\Psi=(V_{\Psi},\mathcal{C}_{\Psi})$ be a $k$-CNF, $\sigma$ be a feasible partial assignment of $\Psi$, and $\Psi^{\sigma}=(V_{\Psi}^{\sigma},\mathcal{C}_{\Psi}^{\sigma})$ be the reduced CNF (see 2.4). Suppose we have
[$\mathbb{P}[\lnot C\mid\sigma]\leq\zeta\qquad\text{for every $C\in\mathcal{C}_{\Psi}^{\sigma}$}.$
@@
We remark that in Algorithm 6, our partial assignment $\sigma_{\Lambda}$ is always supported on the marked variables. Since every clause has at least $\alpha k$ unmarked variables that are not assigned by $\sigma_{\Lambda}$, we get that [$\mathbb{P}[\lnot C\mid\sigma_{\Lambda}]\leq 2^{-\alpha k}$ for every $C$. Thus when applying Proposition B.2, we will take $\zeta=2^{-\alpha k}$.
We can now give an upper bound on the expected cost of Algorithm 6, proving Lemma 4.7.
Proof of Lemma 4.7.
[…]
@@
We now estimate the cost of Line 19. Observe that in Line 19, we can apply the Bernoulli factory in Proposition B.2 to compute the padding distribution $(\nu_{\Psi})_{u}^{\mathrm{pad},\sigma_{\Lambda}}$. Since the connected component $\Psi$ in Line 19 has at most $d|V^{\prime}|\leq 2d^{2}|R|/\alpha$ clauses, by Proposition B.2, the expected cost of Line 19 is at most
@@
Let $R_{\max}=80dk^{4}\log n$. Combining the above and applying 5, we get that the expected cost of $\mathsf{Glauber}(t)$ is at most
@@
$2^{-\frac{1}{48dk^{4}}}\cdot e^{\frac{2d^{2}/\alpha}{2^{(1-\alpha)k}}}\leq 0.9\quad\Rightarrow\quad\sum_{r=0}^{R_{\max}}\left(2^{-\frac{1}{48dk^{4}}}\cdot e^{\frac{2d^{2}/\alpha}{2^{\alpha k}}}\right)^{r}=O(1).$
Appendix C Proof of Theorem 3.3
In this section we show Theorem 3.3. Let $\Phi=(V,\mathcal{C})$ be a $(k,d)$-formula, $\alpha>0$, and $\mathcal{M}\subseteq V$ be an $\alpha$-marking. We introduce the local access algorithm $\mathsf{Conn}$ (Algorithm 7) such that for every unmarked variable $v\in V\setminus\mathcal{M}$, $\mathsf{Conn}(v)$ returns the connected component of $\Phi^{(\mathsf{MarginSample}(u))_{u\in\mathcal{M}}}$ that contains $v$.
@@
12:              Push $\{C^{\prime}\in\mathcal{C}\mid w\in\operatorname{vbl}((C^{\prime})^{\sigma}),(C^{\prime})^{\sigma}\neq 1,C^{\prime}\text{ has never been added to $\mathcal{S}$}\}$ onto $\mathcal{S}$               return $\Phi_{v}^{\sigma}$
@@
To show Theorem 3.3, it remains to show that the algorithm $\mathsf{Conn}(v)$ has expected cost $\mathrm{poly}\log(n)$. We show this by giving an upper bound on the size of the connected component of $\Phi^{(\mathsf{MarginSample}(v))_{u\in\mathcal{M}}}$ that contains $v$.
Proposition C.1.
@@
Recall the dependency hypergraph $H_{\Phi}=(V,\mathcal{E})$ defined in 2.2. Recall also from Definition A.3 that $Z\subseteq V$ is a 2-tree if both of the following conditions are satisfied:
@@
Finally, we union bound over all 2-trees of size $\left\lfloor\frac{b}{kd}\right\rfloor$ that contains $v$. Recall from A.4 that, since $H_{\Phi}$ is a hypergraph with maximum degree $d$, the number of 2-trees $Z\subseteq H_{\Phi}$ containing $v$ with $|Z|=\ell$ is at most $\frac{(ed^{2})^{\ell-1}}{2}$. Setting $\ell=\left\lfloor\frac{b}{kd}\right\rfloor$ and applying a union bound, we get that
@@
Observe that in $\mathsf{Conn}(\cdot)$, every clause is added to $\mathcal{S}$ because its variable set intersects with that of some unsatisfied clause in $\Phi_{v}^{\tau}$. Thus by Proposition C.1, we know that with probability $1-n^{-0.1\log n}$ the number of clauses ever added to $\mathcal{S}$ is at most $\Delta kd\log^{2}n\leq k^{2}d^{2}\log^{2}n$. In this case, Line 7 of Algorithm 7 is executed at most $k^{2}d^{2}\log^{2}n\cdot k=k^{3}d^{2}\log^{2}n$ times. Combined with Lemma 4.7, we get that the expected cost of $\mathsf{Conn}$ is at most
@@
Thus, Theorem 3.3 is proved.
Appendix D Proof of Theorem 2.10
Let $c>0$ be any constant. Given a $k$-uniform hypergraph $H$ and constants $10<\alpha<\beta_{1}<\beta_{2}<$ 1, we describe a local computation algorithm $\mathcal{A}$ that with probability $1-n^{-c}$ finds a 2-coloring of $H$, such that every edge in $H$ has at least $\alpha k$ vertices in each color. Taking $H=H_{\Phi}$ (as in 2.2), we see that finding an $\alpha$-marking for $\Phi$ is equivalent to finding such a 2-coloring for $H_{\Phi}$. The algorithm $\mathcal{A}$ runs in three phases.
@@
Suppose $1/2<\beta_{1}<\beta_{2}<1-\alpha$ satisfy conditions Theorem 2.10:
@@
[…]
@@
$\mathbb{P}[\text{$W$ is contained in some connected component after Phase 1}]\leq p^{w}.$
@@
which is at most $n^{-2c}$ when $w=c_{1}\log n$ with $c_{1}>0$ sufficiently large. By Markov’s inequality, with probability at least $1-n^{-2c}$, there is no 3-tree of size larger than $c_{1}\log n$. Since every size-$s$ connected component in $H$ contains a 3-tree of size at least $s/(dk)^{3}$, we get that with probability at least $1-n^{-2c}$, there is no connected component of size larger than than $c_{1}(dk)^{3}\log n$ in $H$. ∎
Lemma D.2.
@@
The argument in similar to the proof of Lemma D.1. Let $e$ be any edge, and $v$ be any vertex in $\widetilde{H}_{1}$. Observe that
$\mathbb{P}[\text{$e$ becomes dangerous-2 in Phase 2}]\leq 2\binom{(1-\delta)k}{(\beta_{2}-\delta)k}2^{-(\beta_{2}-\delta)k}\text{ for some $0\leq\delta\leq\beta_{1}$},$
@@
which is at most $N^{2}\leq c_{1}^{2}\log^{2}n$ when $w=c_{2}\log N$ with $c_{2}>0$ sufficiently large. By Markov’s inequality, with probability at least $1-1/(c_{1}^{2}\log^{2}n)\geq 1-1/\log^{2}n$, there is no 3-tree of size larger than $c_{2}\log N\leq c_{2}\log(c_{1}\log n)$. This implies that with probability at least $1-1/\log^{2}n$, there is no connected component of size larger than than $c_{2}(dk)^{3}\log(c_{1}\log n)\leq c_{1}c_{2}(dk)^{3}\log\log n$ in $\widetilde{H}_{1}$. ∎
Lemma D.3.
@@
we know from the Lovász local lemma that with positive probability, none of the $A_{e}$’s occurs. ∎
@@
In Phase 1, whether $v$ is colored or troubled-1 completely depends on its descendants in $\mathcal{T}_{v}$. Thus, to compute the assignment of $v$ in Phase 1, it suffices to run Phase 1 on all vertices in $\mathcal{T}_{v}$. By Lemma D.4, with probability $1-n^{-2c}$, this step takes $\mathrm{polylog}(n)$ time.
If $v$ gets colored in Phase 1, then we return the color directly. Otherwise, if $v$ becomes troubled-1 after Phase 1, then we compute the connected component of $H$ containing $v$ after Phase 1. Recall from Lemma D.1 that with probability $1-n^{-2c}$, all connected components after Phase 1 has size at most $c_{1}\log n$, so this step takes another $\mathrm{polylog}(n)$ time.
Let $H_{v}$ denote this connected component. By Lemma D.2, if we perform $O(\log n/\log\log n)$ parallel instances of Phase 2 on $H_{v}$, then with probability $1-n^{-2c}$, in one of them $H_{v}$ is further reduced to connected components each of size $c_{2}\log\log n$. We then perform Phase 3 to find a valid coloring on the component that contains $v$, which by Lemma D.3 is guaranteed to succeed and finishes in $2^{O(\log\log n)}=\mathrm{polylog}(n)$ time.
@@
Appendix E Proof of Lemma 5.2
We first check conditions in Theorem 2.10. For sufficiently large $k$, we have:
  • $0.7784\cdot\frac{1}{75}<2(1-0.96)<1-$ 0.778,
@@
$\qquad\qquad+\frac{\log\left(\frac{0.96\,-\delta}{1-\delta}\right)}{(1-x)\log(2)}-\frac{(0.96\,-x)\log\left(\frac{0.96\,-\delta}{1-\delta}\right)}{(1-\delta)^{2}\log(2)}-\frac{\frac{1}{1-\delta}-\frac{0.96\,-\delta}{(1-\delta)^{2}}}{\log(2)}\bigg{)}-1$
$<0$
for $0.7780\leq\delta\leq\beta_{1}=$ 0.778.
We now check conditions in 5.1. For sufficiently large $k$, we have:
@@
  • $236ed^{3}k^{4}\cdot 0.6^{\alpha k}=36ek^{4}(2^{3/400}\cdot 0.6^{1/75})^{k}\leq 1/$ 2,

### v3

@@ Appendix B Correctness of Algorithm˜5
Let $\mathcal{T}$ denote the set of 2-trees of $V(G)$ of size $\eta+1$ that intersects with $W^{\prime}$. Since $|W^{\prime}|\leq d$ and $G$ has maximum degree at most $6k^{2}d-1$, by Definition˜B.4, we have
@@
where the last step used the assumption that $\theta\geq 0.4$ and $36ed^{3}k^{4}\cdot 0.6^{\alpha k}\leq 1/2$. This implies that
@@
Setting $\eta=(3+c)\log n$, we get the following correctness statement on Algorithm˜5.
Corollary B.12.
@@
Since we have picked $T$ in Algorithm˜3 sufficiently large so that $d_{\operatorname{\mathsf{TV}}}(X_{0},\nu)\leq n^{-(2+c)}$, we get that the joint output $(\mathsf{MarginSample}(u))_{u\in\mathcal{M}}$ satisfies $d_{\operatorname{\mathsf{TV}}}((\mathsf{MarginSample}(u))_{u\in\mathcal{M}},\nu)\leq n^{-c}$, proving Proposition˜4.4.
Appendix C Efficiency of Algorithm˜5
We now move on to show the efficiency of $\mathsf{Glauber}(t)$ for all $t\leq 0$. Observe that for every $r\geq 48dk^{4}$, by Proposition˜B.5, we have
$\hskip 136.5733pt\mathbb{P}[|R|\geq r]\leq 2^{-\frac{r}{48dk^{4}}}.$ (4)
@@ Appendix C Efficiency of Algorithm˜5
We start by upper bounding the size of the final sets $V^{\prime}$ and $\Lambda$ in Algorithm˜5 in terms of $|R|$.
Lemma C.1.
The $V^{\prime}$ and $\Lambda$ in Line 19 of Algorithm˜5 satisfy
@@
Proof C.2.
By Line 10 of Algorithm˜5, we know that for all $u\in\Lambda$, there exists clause $C\in\mathcal{C}$ such that $u\in\operatorname{vbl}(C)$ and $\operatorname{vbl}(C)\cap V^{\prime}\neq\emptyset$. This shows that $|\Lambda|\leq kd|V^{\prime}|$. Observe that $V^{\prime}$ contains at least $|V^{\prime}|/(\Delta+1)\geq|V^{\prime}|/(2dk)$ clauses with disjoint variable sets. Since every marked variable in the clauses added to $V^{\prime}$ have gone through at least one round of LB-Sample (per Line 5 and Line 12), we get that $|R|\geq\alpha k\cdot|V^{\prime}|/(2dk)=\alpha|V^{\prime}|/(2d)$.
@@
Proposition C.3 ([18, Lemma 3.10 and Appendix A]).
Let $\Psi=(V_{\Psi},\mathcal{C}_{\Psi})$ be a $k$-CNF, $\sigma$ be a feasible partial assignment of $\Psi$, and $\Psi^{\sigma}=(V_{\Psi}^{\sigma},\mathcal{C}_{\Psi}^{\sigma})$ be the reduced CNF (see Section˜2). Suppose we have
$\mathbb{P}[\lnot C\mid\sigma]\leq\zeta\qquad\text{for every $C\in\mathcal{C}_{\Psi}^{\sigma}$}.$
@@
We remark that in Algorithm˜5, our partial assignment $\sigma_{\Lambda}$ is always supported on the marked variables. Since every clause has at least $\alpha k$ unmarked variables that are not assigned by $\sigma_{\Lambda}$, we get that $\mathbb{P}[\lnot C\mid\sigma_{\Lambda}]\leq 2^{-\alpha k}$ for every $C$. Thus when applying Proposition˜C.3, we will take $\zeta=2^{-\alpha k}$.
We can now give an upper bound on the expected cost of Algorithm˜5, proving Lemma˜4.5.
Proof C.4 (Proof of Lemma˜4.5).
Let $A$ be the set of all $t^{\prime}$ such that $\mathsf{Glauber}(t^{\prime})$ is executed at least once as subroutine of $\mathsf{Glauber}(t)$. Since we run $\textsf{LB-Sample}(t^{\prime})$ at the beginning (Line 7) of each round $\mathsf{Glauber}(t^{\prime})$, we know that $|A|\leq|R|$.
@@
We now estimate the cost of Line 19. Observe that in Line 19, we can apply the Bernoulli factory in Proposition˜C.3 to compute the padding distribution $(\nu_{\Psi})_{u}^{\mathrm{pad},\sigma_{\Lambda}}$. Since the connected component $\Psi$ in Line 19 has at most $d|V^{\prime}|\leq 2d^{2}|R|/\alpha$ clauses, by Proposition˜C.3, the expected cost of Line 19 is at most
@@
Let $R_{\max}=80dk^{4}\log n$. Combining the above and applying Equation˜4, we get that the expected cost of $\mathsf{Glauber}(t)$ is at most
@@
$2^{-\frac{1}{48dk^{4}}}\cdot e^{\frac{2d^{2}/\alpha}{2^{(1-\alpha)k}}}\leq 0.9\quad\Longrightarrow\quad\sum_{r=0}^{R_{\max}}\left(2^{-\frac{1}{48dk^{4}}}\cdot e^{\frac{2d^{2}/\alpha}{2^{\alpha k}}}\right)^{r}=O(1).$
Appendix D Proof of Theorem˜3.4
In this section we show Theorem˜3.4. Let $\Phi=(V,\mathcal{C})$ be a $(k,d)$-formula, $\alpha>0$, and $\mathcal{M}\subseteq V$ be an $\alpha$-marking. We introduce the local access algorithm $\mathsf{Conn}$ (Algorithm˜7) such that for every unmarked variable $v\in V\setminus\mathcal{M}$, $\mathsf{Conn}(v)$ returns the connected component of $\Phi^{(\mathsf{MarginSample}(u))_{u\in\mathcal{M}}}$ that contains $v$.
@@
12:     Push $\{C^{\prime}\in\mathcal{C}\mid w\in\operatorname{vbl}((C^{\prime})^{\sigma}),(C^{\prime})^{\sigma}\neq 1,C^{\prime}\text{ has not been in $\mathcal{S}$}\}$ onto $\mathcal{S}$      return $\Phi_{v}^{\sigma}$
@@
To show Theorem˜3.4, it remains to show that the algorithm $\mathsf{Conn}(v)$ has expected cost $\mathrm{poly}\log(n)$. We show this by giving an upper bound on the size of the connected component of $\Phi^{(\mathsf{MarginSample}(v))_{u\in\mathcal{M}}}$ that contains $v$.
Proposition D.1.
@@
Recall the dependency hypergraph $H_{\Phi}=(V,\mathcal{E})$ defined in Section˜2. Recall also from Definition˜B.4 that $Z\subseteq V$ is a 2-tree if both of the following conditions are satisfied:
@@
Finally, we union bound over all 2-trees of size $\left\lfloor\frac{b}{kd}\right\rfloor$ that contains $v$. Recall from Definition˜B.4 that, since $H_{\Phi}$ is a hypergraph with maximum degree $d$, the number of 2-trees $Z\subseteq H_{\Phi}$ containing $v$ with $|Z|=\ell$ is at most $\frac{(ed^{2})^{\ell-1}}{2}$. Setting $\ell=\left\lfloor\frac{b}{kd}\right\rfloor$ and applying a union bound, we get that
@@
Observe that in $\mathsf{Conn}(\cdot)$, every clause is added to $\mathcal{S}$ because its variable set intersects with that of some unsatisfied clause in $\Phi_{v}^{\tau}$. Thus by Proposition˜D.1, we know that with probability $1-n^{-0.1\log n}$ the number of clauses ever added to $\mathcal{S}$ is at most $\Delta kd\log^{2}n\leq k^{2}d^{2}\log^{2}n$. In this case, Line 7 of Algorithm˜7 is executed at most $k^{2}d^{2}\log^{2}n\cdot k=k^{3}d^{2}\log^{2}n$ times. Combined with Lemma˜4.5, we get that the expected cost of $\mathsf{Conn}$ is at most
@@
Thus, Theorem˜3.4 is proved.
Appendix E Proof of Theorem˜2.1
Let $c>0$ be any constant. Given a $k$-uniform hypergraph $H$ and constants $0<\alpha<\beta_{1}<\beta_{2}<1$, we describe a local computation algorithm $\mathcal{A}$ that with probability $1-n^{-c}$ finds a 2-coloring of $H$, such that every edge in $H$ has at least $\alpha k$ vertices in each color. Taking $H=H_{\Phi}$ (as in Section˜2), we see that finding an $\alpha$-marking for $\Phi$ is equivalent to finding such a 2-coloring for $H_{\Phi}$. The algorithm $\mathcal{A}$ runs in three phases.
@@
Suppose $1/2<\beta_{1}<\beta_{2}<1-\alpha$ satisfy conditions Theorem˜2.1:
@@
$\mathbb{P}[\text{$e$ becomes dangerous-1 in Phase 1}]$ $\leq 2\binom{k}{\beta_{1}k}2^{-\beta_{1}k},$
$\mathbb{P}[\text{$e$ becomes unsafe-1 in Phase 1}]$ $\leq 2kd\binom{k}{\beta_{1}k}2^{-\beta_{1}k},$
$\mathbb{P}[\text{$v$ lies in a dangerous-1 or unsafe-1}$ $\text{ edge after Phase 1}]$
@@
[…]
@@
which is at most $n^{-2c}$ when $w=c_{1}\log n$ with $c_{1}>0$ sufficiently large. By Markov’s inequality, with probability at least $1-n^{-2c}$, there is no 3-tree of size larger than $c_{1}\log n$. Since every size-$s$ connected component in $H$ contains a 3-tree of size at least $s/(dk)^{3}$, we get that with probability at least $1-n^{-2c}$, there is no connected component of size larger than than $c_{1}(dk)^{3}\log n$ in $H$.
Lemma E.3.
@@
The argument in similar to the proof of Lemma˜E.1. Let $e$ be any edge, and $v$ be any vertex in $\widetilde{H}_{1}$. Observe that
[…]
@@
which is at most $N^{2}\leq c_{1}^{2}\log^{2}n$ when $w=c_{2}\log N$ with $c_{2}>0$ sufficiently large. By Markov’s inequality, with probability at least $1-1/(c_{1}^{2}\log^{2}n)\geq 1-1/\log^{2}n$, there is no 3-tree of size larger than $c_{2}\log N\leq c_{2}\log(c_{1}\log n)$. This implies that with probability at least $1-1/\log^{2}n$, there is no connected component of size larger than than $c_{2}(dk)^{3}\log(c_{1}\log n)\leq c_{1}c_{2}(dk)^{3}\log\log n$ in $\widetilde{H}_{1}$.
Lemma E.5.
@@
we know from the Lovász local lemma that with positive probability, none of the $A_{e}$’s occurs.
@@
In Phase 1, whether $v$ is colored or troubled-1 completely depends on its descendants in $\mathcal{T}_{v}$. Thus, to compute the assignment of $v$ in Phase 1, it suffices to run Phase 1 on all vertices in $\mathcal{T}_{v}$. By Lemma˜E.7, with probability $1-n^{-2c}$, this step takes $\mathrm{polylog}(n)$ time.
If $v$ gets colored in Phase 1, then we return the color directly. Otherwise, if $v$ becomes troubled-1 after Phase 1, then we compute the connected component of $H$ containing $v$ after Phase 1. Recall from Lemma˜E.1 that with probability $1-n^{-2c}$, all connected components after Phase 1 has size at most $c_{1}\log n$, so this step takes another $\mathrm{polylog}(n)$ time.
Let $H_{v}$ denote this connected component. By Lemma˜E.3, if we perform $O(\log n/\log\log n)$ parallel instances of Phase 2 on $H_{v}$, then with probability $1-n^{-2c}$, in one of them $H_{v}$ is further reduced to connected components each of size $c_{2}\log\log n$. We then perform Phase 3 to find a valid coloring on the component that contains $v$, which by Lemma˜E.5 is guaranteed to succeed and finishes in $2^{O(\log\log n)}=\mathrm{polylog}(n)$ time.
@@
Appendix F Proof of Lemma˜5.2
We first check conditions in Theorem˜2.1. For sufficiently large $k$, we have:
  • $4\cdot\frac{1}{75}<2(1-0.96)<1-0.778$,
@@
$\qquad\qquad+\frac{\log\left(\frac{0.96\,-\delta}{1-\delta}\right)}{(1-x)\log(2)}-\frac{(0.96\,-x)\log\left(\frac{0.96\,-\delta}{1-\delta}\right)}{(1-\delta)^{2}\log(2)}-\frac{\frac{1}{1-\delta}-\frac{0.96\,-\delta}{(1-\delta)^{2}}}{\log(2)}\bigg{)}-1<0$
for $0\leq\delta\leq\beta_{1}=0.778$.
We now check conditions in ˜5.1. For sufficiently large $k$, we have:
@@
  • $36ed^{3}k^{4}\cdot 0.6^{\alpha k}=36ek^{4}(2^{3/400}\cdot 0.6^{1/75})^{k}\leq 1/2$,